find the slope of the line containing the given pair of points. if the line is undefind,state so. please be sure to show all of your work. (-4,3)and (3,3)
Answer by jim_thompson5910(35256)

m = (y2-y1)/(x2-x1)


m = (3-3)/(3-(-4))


m = (3-3)/(3+4)


m = 0/7


m = 0


So the slope of the line that passes through the two points is 0
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------